# Unified Multimodal Brain Decoding via Cross-Subject Soft-ROI Fusion

### v1

@@ Abstract
Multimodal brain decoding aims to reconstruct semantic information that is consistent with visual stimuli from brain activity signals such as fMRI, and then generate readable natural language descriptions. However, multimodal brain decoding still faces key challenges in cross-subject generalization and interpretability. We propose a BrainROI model and achieve leading-level results in brain-captioning evaluation on the NSD dataset. Under the cross-subject setting, compared with recent state-of-the-art methods and representative baselines, metrics such as BLEU-4 and CIDEr show clear improvements. Firstly, to address the heterogeneity of functional brain topology across subjects, we design a new fMRI encoder. We use multi-atlas soft functional parcellations (soft-ROI) as a shared space. We extend the discrete ROI Concatenation strategy in MINDLLM to a voxel-wise gated fusion mechanism (Voxel-gate). We also ensure consistent ROI mapping through global label alignment, which enhances cross-subject transferability. Secondly, to overcome the limitations of manual and black-box prompting methods in stability and transparency, we introduce an interpretable prompt optimization process. In a small-sample closed loop, we use a locally deployed Qwen model to iteratively generate and select human-readable prompts. This process improves the stability of prompt design and preserves an auditable optimization trajectory. Finally, we impose parameterized decoding constraints during inference to further improve the stability and quality of the generated descriptions.
@@ 1 Introduction
Non-invasive neuroimaging (e.g., fMRI) provides observable neural representations for studying human visual cognition. With the development of vision-language large models (VLMs/MLLMs), recent methods attempt to map brain activity into a visual semantic feature space and perform decoding tasks such as image captioning, localization, and question answering on frozen multimodal language models. A representative work, UMBRAE [1], adopts the "fMRI → visual space → frozen MLLM" paradigm, demonstrating the potential of unified zero-shot multi-task decoding while avoiding the cost of end-to-end large-model fine-tuning. However, UMBRAE still leaves room for improvement in cross-subject generalization: to deal with inter-subject variability, it introduces subject-specific tokenizers, which makes the model contain subject-dependent parameters, hindering the realization of a fully parameter-shared, subject-agnostic solution.
To improve cross-subject robustness, MINDLLM [2] explores using neuroscience priors, namely functional brain regions of interest (ROIs), as shared anchors across subjects. This design aims to reduce representation instability caused by the spatial sparsity of fMRI signals and inter-subject variability. However, MINDLLM relies on a discrete feature concatenation strategy. This strategy makes it hard to achieve an adaptive trade-off between coarse ROI priors and fine-grained voxel-level information. As a result, the question remains open: how can we build a more effective unified mechanism that balances a transferable shared space with detailed voxel-level representations? Meanwhile, IPO [3] introduces interpretable prompt optimization for VLM tasks. IPO treats an LLM as an "optimizer". The method iteratively generates human-readable candidate prompts in a small-sample closed loop. The method then selects the best prompt using target-task metrics. This design offers both strong performance and interpretability. However, existing work has not systematically integrated this idea into the fMRI→MLLM brain-decoding pipeline. Finally, the quality of the generated text is also crucial. Brain-captioning benchmarks mainly rely on metrics such as BLEU and CIDEr. If a model’s generation distribution shows systematic bias relative to the references, the scores can drop sharply. Common biases include outputs that are too short or too long, and repeated n-grams. These issues can hurt both automatic metrics and human readability, even when the overall meaning is roughly correct.
@@
    We propose a cross-subject shared space based on multi-atlas soft ROIs, together with global label alignment. We also introduce voxel-wise gated fusion (Voxel-gate). This design balances transferability and fine-grained spatial representation.
@@ 2.2 Cross-subject Generalization
2.3 Prompt Optimization for VLM/MLLM
@@ 3.1 Overall Framework
Our brain-decoding framework follows a three-stage pipeline as illustrated in Fig. 1: First, we design an fMRI encoder based on multi-atlas soft functional parcellations (soft-ROI) with gated fusion. This encoder enables efficient and robust cross-subject feature extraction. Next, we introduce an LLM-based closed-loop optimization process. This process automatically generates and selects optimal human-readable prompts. Finally, we feed the encoded brain-activity features and the optimized prompts into a frozen MLLM with parameterized constraints, and the model generates high-quality descriptive text.
Figure 1: Overall framework of our three-stage brain-decoding pipeline.
@@
Next, we map $R^{(a,s)}$ to an ROI prior embedding $E^{(a)}=R^{(a,s)}W_{a}$, and provide three switchable fusion strategies along the atlas dimension: Concatenation, Gate (voxel-independent global weighting with $\alpha_{a}$), and Voxel-gate (voxel-dependent locally adaptive weighting with $\alpha_{n,a}$). Here, soft-ROI refers to the fused $\tilde{E}_{n}$ (the soft fusion result of multi-atlas priors), rather than probabilistic membership within a single atlas.
@@ 3.2.3 Key Construction
As shown in Fig. 2, in the Gate strategy branch, we first apply average pooling over the voxel dimension to the ROI embeddings from each atlas, yielding an atlas-specific global summary vector of dimension $d_{\mathrm{roi}}$. All atlas summaries are then stacked to form an atlas-level representation of size $A\times d_{\mathrm{roi}}$. Next, a gating MLP produces a score for each atlas summary vector (with shared parameters across atlases), and a softmax over the atlas dimension yields the global weights $\boldsymbol{\alpha}$. These weights are shared across all voxels of the same sample, so the ROI embeddings from multiple atlases can be fused using the same set of weights, resulting in a fused atlas-level ROI representation of size $N_{s}\times d_{\mathrm{roi}}$. Finally, the voxel coordinate encoding $C$ (of size $N_{s}\times d_{c}$) is concatenated with the fused ROI representation along the feature dimension to form $X\in\mathbb{R}^{N_{s}\times(d_{c}+d_{\mathrm{roi}})}$, which is then linearly projected to obtain the keys $K$. This strategy offers good interpretability and low computational overhead; however, since $\boldsymbol{\alpha}$ is shared across voxels, its ability to adapt to voxel-wise local variations is limited.
@@ 4.1 Experimental Setup
We conduct all experiments on the Natural Scenes Dataset (NSD). A brief overview of NSD is provided in Appendix 0.A. We strictly follow the standard protocol used in visual decoding studies to ensure reproducibility and fair comparison.
@@ 4.3.1 Ablation on Cross-subject Fusion Strategy
Table 2: Ablation results of cross-subject fusion strategies.
@@
As the result shown in Table 2. Voxel-wise gating (Voxel-gate) achieves the best performance on 7 out of 8 metrics. The only exception is RefCLIP-S, where Gate is higher by a very small margin (0.0004). The advantage of Voxel-gate is more pronounced on metrics that are sensitive to higher-order $n$-gram structure and content coverage. Compared with Concatenation, Voxel-gate improves BLEU-4 from 0.2837 to 0.2911 and improves CIDEr from 0.6629 to 0.6952. It also yields consistent gains on BLEU-1/2/3, ROUGE-L, CLIP-S, and RefCLIP-S. Compared with Gate, Voxel-gate improves BLEU-4 by 0.0022 and CIDEr by 0.0097, while BLEU-1/2/3 and ROUGE-L also increase. We note that CLIP-S and RefCLIP-S are nearly tied across the three fusion modes. This suggests that their global semantic similarity is comparable. In contrast, Voxel-gate shows a stable advantage on BLEU-3/4 and CIDEr, which are more sensitive to fine-grained structure and content consistency. This result indicates that Voxel-gate is more effective at recovering syntactic patterns and detailed semantics.
@@ 4.3.2 Ablation on Constrained Decoding Strategy
Table 3: Ablation settings for constrained decoding.
@@
Table 4: Ablation results for constrained decoding strategies.
@@
4.3.3 Subject-specific Evaluation under Cross-subject Joint Training
@@ Appendix 0.A Natural Scenes Dataset (NSD) Overview
The Natural Scenes Dataset (NSD) is a large-scale natural-scene fMRI benchmark for visual neuroscience and NeuroAI. NSD was acquired with whole-brain ultra-high-field 7T fMRI (1.8 mm isotropic resolution; TR = 1.6 s) and, through a multi-session longitudinal design, provides high-volume, high-SNR visually evoked responses for each individual subject. Data collection typically spans 30–40 scan sessions (distributed weekly over roughly one year). Each subject views approximately 9,000–10,000 distinct color natural-scene images, totaling about 22,500–30,000 stimulus presentations. The stimuli are mainly drawn from natural image repositories such as MS COCO, covering diverse objects, scenes, and compositional relationships, which offer substantial semantic diversity and difficulty for tasks such as predicting visual semantics or descriptive text from brain signals.
@@
Appendix 0.B Interpretable Prompt Optimization (IPO) and an Example Trace
@@
Table 6: Example IPO trace ranked by validation BLEU-4.
@@ Appendix 0.C Qualitative Examples of Brain Captioning
Figure 4: Qualitative reference–candidate caption comparisons.

### v2

@@ Abstract
Multimodal brain decoding aims to reconstruct semantic information that is consistent with visual stimuli from brain activity signals such as fMRI, and then generate readable natural language descriptions. However, multimodal brain decoding still faces key challenges in cross-subject generalization and interpretability. We propose a BrainROI model and achieve leading-level results in brain-captioning evaluation on the Natural Scenes Dataset (NSD) dataset. Under the cross-subject setting, compared with recent state-of-the-art methods and representative baselines, metrics such as BLEU-4 and CIDEr show clear improvements. Firstly, to address the heterogeneity of functional brain topology across subjects, we design a new fMRI encoder. We use multi-atlas soft functional parcellations (soft-ROI) as a shared space. We extend the discrete Regions of Interest (ROI) Concatenation strategy in MINDLLM to a voxel-wise gated fusion mechanism (Voxel-gate). We also ensure consistent ROI mapping through global label alignment, which enhances cross-subject transferability. Secondly, to overcome the limitations of manual and black-box prompting methods in stability and transparency, we introduce an interpretable prompt optimization process. In a small-sample closed loop, we use a locally deployed Qwen model to iteratively generate and select human-readable prompts. This process improves the stability of prompt design and preserves an auditable optimization trajectory. Finally, we impose parameterized decoding constraints during inference to further improve the stability and quality of the generated descriptions.
@@ 1 Introduction
Non-invasive neuroimaging (e.g., fMRI) provides observable neural representations for studying human visual cognition. With the development of vision-language large models (VLMs/MLLMs), recent methods attempt to map brain activity into a visual semantic feature space and perform decoding tasks such as image captioning, localization, and question answering on frozen multimodal language models. A representative work, UMBRAE [1], adopts the "fMRI → visual space → frozen MLLM" paradigm, demonstrating the potential of unified zero-shot multi-task decoding while avoiding the cost of end-to-end large-model fine-tuning. However, UMBRAE still leaves room for improvement in cross-subject generalization: to deal with inter-subject variability, it introduces subject-specific tokenizers, which make the model contain subject-dependent parameters, hindering the realization of a fully parameter-shared, subject-agnostic solution.
To improve cross-subject robustness, MINDLLM [2] explores using neuroscience priors, namely functional brain regions of interest (ROIs), as shared anchors across subjects. This design aims to reduce representation instability caused by the spatial sparsity of fMRI signals and inter-subject variability. However, MINDLLM relies on a discrete feature concatenation strategy. This strategy makes it hard to achieve an adaptive trade-off between coarse ROI priors and fine-grained voxel-level information. As a result, the question remains open: how can we build a more effective unified mechanism that balances a transferable shared space with detailed voxel-level representations? Meanwhile, IPO [3] introduces interpretable prompt optimization for MLLM tasks. IPO treats an LLM as an "optimizer". The method iteratively generates human-readable candidate prompts in a small-sample closed loop. The method then selects the best prompt using target-task metrics. This design offers both strong performance and interpretability. However, existing work has not systematically integrated this idea into the fMRI→MLLM brain-decoding pipeline. Finally, the quality of the generated text is also crucial. Brain-captioning benchmarks mainly rely on metrics such as BLEU and CIDEr. If a model’s generation distribution shows systematic bias relative to the references, the scores can drop sharply. Common biases include outputs that are too short or too long, and repeated n-grams. These issues can hurt both automatic metrics and human readability, even when the overall meaning is roughly correct.
@@
    We propose a cross-subject shared space based on multi-atlas soft ROIs, together with global label alignment. We also introduce Voxel-gate. This design balances transferability and fine-grained spatial representation.
@@ 2.2 Cross-subject Generalization
2.3 Prompt Optimization for MLLM
@@ 3.1 Overall Framework
Our brain-decoding framework follows a three-stage pipeline as illustrated in Fig. 1: First, we design an fMRI encoder based on multi-atlas (soft-ROI) with gated fusion. This encoder enables efficient and robust cross-subject feature extraction. Next, we introduce an LLM-based closed-loop optimization process. This process automatically generates and selects optimal human-readable prompts. Finally, we feed the encoded brain-activity features and the optimized prompts into a frozen MLLM with parameterized constraints, and the model generates high-quality descriptive text.
Figure 1: Overall framework of our three-stage brain-decoding pipeline. The framework illustrates the integration of soft-ROI fusion, Voxel-gate mechanism, and interpretable prompt optimization.
@@
Next, we map $R^{(a,s)}$ to an ROI prior embedding $E^{(a)}=R^{(a,s)}W_{a}$, and provide three switchable fusion strategies along the atlas dimension: Concatenation, Gate (voxel-independent global weighting with $\alpha_{a}$), and Voxel-gate (with $\alpha_{n,a}$).
@@ 3.2.3 Key Construction
As shown in Fig. 2, in the Gate strategy branch, we first apply average pooling over the voxel dimension to the ROI embeddings from each atlas, yielding an atlas-specific global summary vector of dimension $d_{\mathrm{roi}}$. All atlas summaries are then stacked to form an atlas-level representation of size $A\times d_{\mathrm{roi}}$. Next, a gating multi-layer perceptron produces a score for each atlas summary vector (with shared parameters across atlases), and a softmax over the atlas dimension yields the global weights $\boldsymbol{\alpha}$. These weights are shared across all voxels of the same sample, so the ROI embeddings from multiple atlases can be fused using the same set of weights, resulting in a fused atlas-level ROI representation of size $N_{s}\times d_{\mathrm{roi}}$. Finally, the voxel coordinate encoding $C$ (of size $N_{s}\times d_{c}$) is concatenated with the fused ROI representation along the feature dimension to form $X\in\mathbb{R}^{N_{s}\times(d_{c}+d_{\mathrm{roi}})}$, which is then linearly projected to obtain the keys $K$. This strategy offers good interpretability and low computational overhead; however, since $\boldsymbol{\alpha}$ is shared across voxels, its ability to adapt to voxel-wise local variations is limited.
@@ 4.1 Experimental Setup
We conduct all experiments on the NSD. A brief overview of NSD is provided in Appendix 0.A. We strictly follow the standard protocol used in visual decoding studies to ensure reproducibility and fair comparison.
@@ 4.3.1 Ablation on Cross-subject Fusion Strategy
Table 2: Ablation results of cross-subject fusion strategies, including Concatenation, Gate, and Voxel-gate.
@@
As the results shown in Table 2. Voxel-gate achieves the best performance on 7 out of 8 metrics. The only exception is RefCLIP-S, where Gate is higher by a very small margin (0.0004). The advantage of Voxel-gate is more pronounced on metrics that are sensitive to higher-order $n$-gram structure and content coverage. Compared with Concatenation, Voxel-gate improves BLEU-4 from 0.2837 to 0.2911 and improves CIDEr from 0.6629 to 0.6952. It also yields consistent gains on BLEU-1/2/3, ROUGE-L, CLIP-S, and RefCLIP-S. Compared with Gate, Voxel-gate improves BLEU-4 by 0.0022 and CIDEr by 0.0097, while BLEU-1/2/3 and ROUGE-L also increase. We note that CLIP-S and RefCLIP-S are nearly tied across the three fusion modes. This suggests that their global semantic similarity is comparable. In contrast, Voxel-gate shows a stable advantage on BLEU-3/4 and CIDEr, which are more sensitive to fine-grained structure and content consistency. These results indicate that Voxel-gate is more effective at recovering syntactic patterns and detailed semantics.
@@ 4.3.2 Ablation on Constrained Decoding Strategy
Table 3: Ablation settings for constrained decoding, focusing on repetition suppression and length-preference adjustment.
@@
Table 4: Ablation results for constrained decoding strategies, including repetition suppression and length-preference adjustment.
@@
4.3.3 Subject-specific evaluation under cross-subject joint training (S1, S2, S5, S7).
@@ Appendix 0.A Natural Scenes Dataset (NSD) Overview
The NSD is a large-scale natural-scene fMRI benchmark for visual neuroscience and NeuroAI. NSD was acquired with whole-brain ultra-high-field 7T fMRI (1.8 mm isotropic resolution; TR = 1.6 s) and, through a multi-session longitudinal design, provides high-volume, high-SNR visually evoked responses for each individual subject. Data collection typically spans 30–40 scan sessions (distributed weekly over roughly one year). Each subject views approximately 9,000–10,000 distinct color natural-scene images, totaling about 22,500–30,000 stimulus presentations. The stimuli are mainly drawn from natural image repositories such as MS COCO, covering diverse objects, scenes, and compositional relationships, which offer substantial semantic diversity and difficulty for tasks such as predicting visual semantics or descriptive text from brain signals.
@@
Appendix 0.B IPO and an Example Trace
@@
Table 6: Example IPO trace ranked by validation BLEU-4 score.
@@ Appendix 0.C Qualitative Examples of Brain Captioning
Figure 4: Qualitative comparison of COCO reference captions and fMRI-generated candidate captions.